OpenStudy (anonymous):

Convert 5.764764764... to a rational expression in the form of a over b where b ≠ 0. 5 and 764 over 999 5764 over 9999 5 and 764 over 99 5 and 999 over 764

OpenStudy (rational):

\[5\cdot\overline{764}=5+0.\overline{764}=5+\dfrac{764}{999}=5\dfrac{764}{999}\]